Presented below are selected transactions for Bridgeport Company during September and October of the current year. Bridgeport uses a perpetual inventory system.

Sept. 1 Purchased merchandise on account from Hillary Company at a cost of $45,000, FOB destination, terms 1/15, n/30.
2 The correct company paid $2,000 of freight charges to Trucking Company on the September 1 merchandise purchase.
5 Returned for credit $3,000 of damaged goods purchased from Hillary Company on September 1.
15 Sold the remaining merchandise purchased from Hillary Company to Irvine Company for $70,000, terms 2/10, n/30, FOB destination.
16 The correct company paid $2,200 of freight charges on the September 15 sale of merchandise.
17 Issued Irvine Company a credit of $5,900 for returned goods. These goods had cost Bridgeport Company $3,000 and were returned to inventory.
25 Received the balance owing from Irvine Company for the September 15 sale.
30 Paid Hillary Company the balance owing for the September 1 purchase.
Oct. 1 Purchased merchandise on account from Kimmel Company at a cost of $52,000, terms 2/10, n/30, FOB shipping point.
2 The correct company paid freight costs of $1,100 on the October 1 purchase.
3 Obtained a purchase allowance of $2,000 from Kimmel Company to compensate for some minor damage to goods purchased on October 1.
10 Paid Kimmel Company the amount owing on the October 1 purchase.
11 Sold all of the merchandise purchased from Kimmel Company to Kieso Company for $90,000, terms 2/10, n/30, FOB shipping point.
12 The correct company paid $800 freight costs on the October 11 sale.
17 Issued Kieso Company a sales allowance of $2,000 because some of the goods did not meet Kieso's exact specifications.
31 Received a cheque from Kieso Company for the balance owing on the October 11 sale.

Prepare journal entries to record the above transactions for Bridgeport Company. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Record journal entries in the order presented in the problem.)

~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

At the beginning of the current tennis season, on April 1, 2017, Novak Tennis Shop's inventory consisted of 50 tennis racquets at a cost of $50 each. Novak uses a perpetual inventory system. The following transactions occurred in April:

Apr. 2 Purchased 160 additional racquets from Roberts Inc. for $8,000, terms n/30.
4 Determined that five of the racquets purchased on April 2 were damaged and returned them to Roberts Inc. Roberts Inc. credited Novak’s account.
5 Sold 45 racquets to Tennis Dome for $130 each, terms n/30.
6 Tennis Dome returned 15 of the racquets after determining it had purchased more racquets than it needed. Novak gave Tennis Dome a credit on its account and returned the racquets to inventory.
10 Sold 40 racquets at $130 each to cash customers.
12 Ten of these racquets were returned for cash. The customers claimed they never play tennis and had no idea how they had been talked into purchasing the racquets. Refunded cash to these customers and returned the racquets to inventory.
17 An additional 10 of the racquets sold on April 10 were returned because the racquets were damaged. The customers were refunded cash and the racquets were sent to a local recycler.
25 Sold 60 racquets to the Summer Club for $130 each, terms n/30.
29 Summer Club returned 25 of the racquets after the tennis pro had examined them and determined that these racquets were of inferior quality. Novak gave Summer Club a credit and decided to return the racquets to inventory with plans to sell them for the reduced price of $75 each.

Record the transactions for the month of April for Novak. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Record journal entries in the order presented in the problem.)
Answer #1
Bridgeport Company
Journal entries
Date Account Debit Credit Calculation
1-Sep Inventory $     45,000
Accounts Payable $     45,000
2-Sep No entry
In FOB destination sale, freight is paid by seller
5-Sep Accounts Payable $        3,000
Inventory $        3,000
15-Sep Accounts Receivable $     70,000
Sales Revenue $     70,000
Cost of goods sold $     42,000 =45,000-3,000
Inventory $     42,000
16-Sep Freight charges $        2,200
Cash $        2,200
17-Sep Sales Returns and Allowances $        5,900
Accounts Receivable $        5,900
Inventory $        3,000
Cost of goods sold $        3,000
25-Sep Cash $     62,818
Sales Discounts $        1,282 =64,100*2%
Accounts Receivable $     64,100 =70,000-5,900
30-Sep Accounts Payable $     42,000 =45,000-3,000
Cash $     42,000
Oct
1-Oct Inventory $     52,000
Accounts Payable $     52,000
2-Oct Inventory $        1,100
Cash $        1,100
3-Oct Accounts Payable $        2,000
Inventory $        2,000
10-Oct Accounts Payable $     50,000 =52,000-2,000
Cash $     49,000
Inventory $        1,000 =50,000*2%
11-Oct Accounts Receivable $     90,000
Sales Revenue $     90,000
Cost of goods sold $     50,100 =52,000+1,100-2,000-1,000
Inventory $     50,100
12-Oct No entry
In FOB shipping sale, freight is paid by buyer
17-Oct Sales Returns and Allowances $        2,000
Accounts Receivable $        2,000
31-Oct Cash $     88,000 =90,000-2,000
Accounts Receivable $     88,000

As HOMEWORKLIB, we are required to do 1 question only.